# Fast and Globally Consistent Normal Orientation based on the Winding Number Normal Consistency

### v1

@@
Abstract
Estimating a consistently oriented normal vector field for an unoriented point cloud enables a number of important downstream applications in computer graphics. While normal estimation for a small patch of points can be done with simple techniques like principal component analysis (PCA), orienting these normals to be globally consistent has been a notoriously difficult problem. Some recent methods exploit various properties of the winding number formula to achieve global consistency with state-of-the-art performance. Despite their exciting progress, these algorithms either have high space/time complexity, or do not produce accurate and consistently oriented normals for imperfect data. In this paper, we derive a novel property from the winding number formula to tackle this problem: the normal consistency property of the winding number formula. We refer to this property as the winding number normal consistency (WNNC). The derived property is based on the simple observation that the normals (negative gradients) sampled from the winding number field should be codirectional to the normals used to compute the winding number field. We further propose to turn the WNNC property into a normal update formula, which leads to an embarrassingly simple yet effective iterative algorithm that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, our proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives, which can be accelerated and parallelized using treecode-based approximation algorithms due to their special structures. Exploiting this fact, we implement a GPU-accelerated treecode-based solver. Our GPU (and even CPU) implementation can be significantly faster than the recent state-of-the-art methods for normal orientation from raw points.
@@ Abstract
Figure 1: Given a raw point cloud, our method iteratively updates its normals to be globally consistently oriented (points rendered as spheres colored by normals). Our iterative algorithm produces overall consistent normals even at the first iteration, and fully converges in around 40 iterations. With our treecode-accelerated implementation running on GPU (RTX 3090), the algorithm obtains high-fidelity normals in $\sim$1 second for 50,000 points, and $\sim$30 seconds for a point cloud with half a million points.
1 Introduction
@@
Point clouds are one of most widely used 3D representations in computer graphics applications such as geometric modeling and rendering. In most cases, consistently oriented normal vectors associated to points are necessary, e.g., for surface reconstruction [24, 23, 25, 32]. While locally estimating normals for a small patch of points can be done with simple techniques such as principal component analysis (PCA), obtaining a globally consistent orientation remains a notoriously difficult problem. Older methods mostly take a propagation-based approach by first estimating normals locally and then propagating the orientation globally [18, 42, 27, 37, 34]. These methods are generally fast but less robust to noise, thin structures and sharp edges where propagation easily fails. Recent developments focus more on the global consistency by introducing new optimization formulations that involve the whole point cloud [19, 31, 44]. Despite achieving the state-of-the-art performance, these global methods often suffer from high space/time complexity, preventing their practicability.
@@
An exciting and inspiring tread in this field is the exploration of various properties of the winding number formula (see Theorem 6). From the second line in Eq. (5) of the winding number formula, Lin et al. [31] observed that the winding number field computed using correct normals should evaluate to $1/2$ at the input points. Lin et al. [31] further turned this observation into a practical linear system, from which normals can be solved using conjugate gradients [17]. Complementarily, Xu et al. [44] use the first and the third line in Eq. (5) to define an regularization term such that the winding number evaluates to either 00 or $1$ almost everywhere. By turning the winding number formula to optimization objectives, these two methods above successfully achieve state-of-the-art performance in estimating consistently oriented normals. However, both methods suffer from high computational complexity. The linear system in PGR [31] has size $N\times N$ and is dense. Running PGR [31] with 40,000 points consumes about 8GB GPU memory. The high memory usage prohibits it from scaling to large point clouds. On the other hand, GCNO [44] has high time complexity due to its objective function being non-linear. Solving a 10,000-point model using GCNO takes about one hour.
@@
While it seems that PGR [31] and GCNO [44] have exhausted all possibilities with the winding number formula by using all three lines in Eq. (5) as constraints, we further exploit the winding number formula by exploring its higher-order properties. To be more specific, we propose to utilize the fact that, when given correctly oriented normals to evaluate the winding number formula, the normals (negative gradients) of the formula should be codirectional to the given normals. We refer to this property as the winding number normal consistency (WNNC). An illustration of WNNC is given in Fig. 2. Based on the WNNC, we design an iterative algorithm for updating normals that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, the proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives. Due to their special structures, these operations can be accelerated using a treecode-based algorithm [4] and parallelized on GPU. Lacking an open-source implementation of the treecode algorithm that fully supports our needs (GPU-based and support for the winding number kernel derivative functions), we implement our custom treecode-based solver in CUDA. As shown in Table 1, our implementation significantly outperforms recent state-of-the-art methods in terms of space/time complexity.
@@
  • We implement a treecode-based acceleration algorithm in CUDA, achieving substantial improvements over prior methods in terms of computational costs.
Figure 2: Illustration of the winding number normal consistency (WNNC) property. From left to right, we start with an input point cloud with normals, evaluate the winding number formula, and finally compute the negative field gradients at the input point positions. If the normals at the start are correctly oriented as shown in the first row, then the intermediate field is exactly the winding number field of the underlying shape. Consequently, the negative gradients at these point positions would be aligned with the input normals. We refer to this property as the winding number normal consistency (WNNC) property. Note that random normals do not satisfy the WNNC property (second row).
Table 1: A brief comparison of time and memory consumption for the state-of-the-art methods GCNO [44], PGR [31], Dipole [34] and iPSR [19]. The results are obtained from running an Armadillo model [28] with uniform sampling on a desktop with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU. Note that the cells colored in gray are estimates based on complexity, because GCNO and PGR are too costly to run on large-scale point clouds.
Method $10^{4}$ points $10^{5}$ points $10^{6}$ points
Time Mem. Time Mem. Time Mem.
GCNO (CPU) 1.1h $<$1G $>$1day 1.7G weeks $>$10G
PGR (GPU) 8s 1G $>$10min $>$30G $>$1day $>$3000G
Dipole (GPU) 17s 1G 90s 2.3G 230s 7.5G
iPSR (CPU) 18s $<$1G 50s $<$1G 380s 2.6G
Ours (CPU) 3s $<$1G 15s $<$1G 294s 1G
Ours (GPU) 2s $<$1G 6s $<$1G 86s 1.5G
2 Related Work
The research into estimating consistent normal orientation has a long history, dating back to decades ago. In this review, we mainly introduce methods that focus on the global consistency of normal orientation. Following prior work, we categorize existing methods into three types: propagation-based, volumetric and deep learning-based. We remark that these types are not completely mutually exclusive, and we introduce existing work based on their most prominent contribution.
2.1 Propagation-based Methods
@@
Local normal estimation is usually carried out by fitting a primitive geometry to local samples. Popular choices for this step include local tangent planes obtained from Principal Component Analysis (PCA) [18], implicit quadric surfaces [42, 43], polynomials obtained by moving least squares [29], osculating jets [6, 34]. These techniques are often efficient and well-established.
Once unoriented normals are available, it remains to determine a sign for each of them. The general practice is to propagate local orientations through neighbors to achieve global consistency. Hoppe et al. [18] build a graph for input points whose edge costs are given by a measure of orientation consistency. Seeking a globally consistent orientation is then formulated as a graph optimization problem. Hoppe et al. [18] use the minimum spanning tree (MST) as a greedy solution, which boils down to propagating the orientation along the MST. This simple greedy approach is efficient but lacks robustness. Many follow-up methods improve one or more aspects of Hoppe et al. [18]. To be robust to sharp features, Xie et al. [42] start from multiple seed points with an improved flipping criterion to orient the whole point cloud. König and Gumhold [27] further introduced Hermite curves into the flipping criterion to reduce the possibility of failures. Seversky et al. [37] use harmonic functions on the point cloud to improve the constructed MST. Going beyond MST-based approaches, other propagation schemes have been proposed as well. Jakob et al. [22] build a graph of points and equate a graph-edge collapse propagation with as an orientation step, and implement a fast parallel algorithm that collapse edges in a greedy manner. Instead of propagating orientations on the surface, Xie et al. [43] first use locally oriented surfaces to determine locally oriented regions, and then propagate region orientations through a voting scheme. Recently, Metzer et al. [34] propose to use neural networks to estimate consistently orientated normals for local patches, and then use dipole-based propagation to achieve global consistency.
@@
2.2 Volumetric Methods
Instead of directly estimating normals for point clouds, volumetric methods adopt an indirect approach where the whole volume is partitioned into inside/outside regions. Normal orientations can be then be determined as the direction pointing from inside to outside.
Early works make use of space partitioning structures to identify inside/outside regions. Dey and Goswami [10] prove that, under certain assumptions, Delaunay balls of the Voronoi diagram can well separated into outer balls and inner balls to give a desired partitioning. Chen et al. [7] uses an octree, called binary orientation tree, that gradually tags its node corners and inside/outside as it grows.
@@ 2.2 Volumetric Methods
For applying additional constraints, Alliez et al. [1] make use of unoriented normals estimated from Voronoi cells, and solve for an implicit field that minimizes an anisotropic Dirichlet energy measuring the unoriented alignment between Voronoi-based normals and the implicit field gradient. Walder et al. [40] apply value and gradient maximization together with smoothness regularization to an implicit function represented by radial basis functions (RBFs). Also based on RBFs, Huang et al. [20] adopt triharmonic kernels and Duchon’s energy [11] and reported high robustness to point sparsity and non-uniform sampling.
@@
Signed distance fields (SDFs) are a popular choice of implicit fields. Mello et al. [33] compute an approximate SDF on a tetrahedral grid. Zhao et al. [46] solves the Eikonal equation to obtain the SDF on a rectangular grid. Gropp et al. [15] turns the Eikonal equation as a loss function supervising a neural network representation of the SDF.
@@
Most closely related to our work are methods based on the winding number field [3, 32]. Simply speaking, the winding number field of an object equals $1$ on the inside and 00 on the outside. Since computing the winding number field requires known normals, several methods derive properties of the winding number field and use them in turn as constraints to obtain normals. Observing the fact that the winding number field is piecewise constant (i.e., having zero Dirichlet energy), Takayama et al. [38] solve for an orientation that minimizes the Dirichlet energy of the field computed with the winding number formula. Recently, Lin et al. [31] derived a linear system formulation enforcing the field to equal $1/2$ on input samples, from which normals can be efficiently solved. Xu et al. [44] designed a double-well loss function encouraging the field to evaluate to either 00 or $1$ almost everywhere.
@@
Compared with propagation-based methods, most volumetric methods take a more “global” view, in the sense that their optimization objectives considers all input points throughout¸ the whole algorithm execution process. As a result, these methods generally achieve much better global consistency but at a notably higher computational cost.
2.3 Deep Learning Methods
@@
In the past decade, deep learning has stormed over nearly all scientific fields, with normal estimation being no exception. The powerful fitting ability of neural networks allows direct inference of normals given an input point cloud. PCPNet [16] and Nesti-Net [5] use neural networks to process multi-scale point features and directly infer point normals. There are also volumetric learning-based methods that infer the implicit rather than the normals [12, 41]. Different from these straight-forward regression approaches, AdaFit [48] predicts point offsets prior to normal estimation to reduce the number of outliers. NeAF [30] further extends this idea by learning a neural angle field that offsets any given normal to the ground-truth one.
@@
2.4 Other Consistent Normal Estimation Methods
A few methods do not belong to the aforementioned categories. Shape as Points (SAP) [36] derives a differentiable version of Poisson Surface Reconstruction (PSR) [24], and uses it to update normals such that the reconstructed mesh has a low error with the input points. Since the differentiable PSR [36] is performed in the frequency domain with fast Fourier transform, its reconstruction accuracy is limited by grid resolution. Iterative Poisson Surface Reconstruction (iPSR) [19] starts with applying Screened Poisson Surface Reconstruction (SPSR) [23] to points with arbitrary normals, and iteratively resample normals from the mesh reconstructed in the previous iteration. Hou et al. [19] show that this simple iterative procedure can update normals into a consistent orientation in only 5$\sim$30 of steps.
3 Preliminaries
Our method builds on the winding number formula [3] (also referred to as the Gauss formula in potential theory [32]), and is closely related to PGR [31]. We begin with a brief review of them.
Theorem 1 (Winding Number).

### v2

@@
Abstract.
Estimating consistently oriented normals for point clouds enables a number of important applications in computer graphics such as surface reconstruction. While local normal estimation is possible with simple techniques like principal component analysis (PCA), orienting these normals to be globally consistent has been a notoriously difficult problem. Some recent methods exploit various properties of the winding number formula to achieve global consistency with state-of-the-art performance. Despite their exciting progress, these algorithms either have high space/time complexity, or do not produce accurate and consistently oriented normals for imperfect data. In this paper, we propose a novel property from the winding number formula, termed Winding Number Normal Consistency (WNNC), to tackle this problem. The derived property is based on the simple observation that the normals (negative gradients) sampled from the winding number field should be codirectional to the normals used to compute the winding number field. Since the WNNC property itself does not resolve the inside/outside orientation ambiguity, we further propose to incorporate an objective function from Parametric Gauss Reconstruction (PGR). We propose to iteratively update normals by alternating between WNNC-based normal updates and PGR-based gradient descents, which leads to an embarrassingly simple yet effective iterative algorithm that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, our proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives, which can be accelerated and parallelized using a treecode-based approximation algorithm due to their special structures. Exploiting this fact, we implement a GPU-accelerated treecode-based solver. Our GPU (and even CPU) implementation can be significantly faster than the recent state-of-the-art methods for normal orientation from raw points. Our code is integrated with the popular PyTorch framework to facilitate further research into winding numbers, and is publicly available at https://jsnln.github.io/wnnc/index.html.
Normal orientation, raw point cloud, winding number
copyright: noneccs: Computing methodologies Shape modeling
@@ Abstract.
Figure 1. Given a raw point cloud, our method iteratively updates its normals to be globally consistently oriented (points are rendered as spheres colored by normals). Our iterative algorithm produces overall consistent normals even at the first iteration, and fully converges in around 40 iterations. With our treecode-accelerated implementation running on GPU (RTX 3090), the algorithm obtains high-fidelity normals in $\sim$1 second for 50,000 points, and $\sim$30 seconds for a point cloud with half a million points.
[…]
Teaser.
1. Introduction
@@
Point clouds are one of most widely used 3D representations in computer graphics applications such as geometric modeling and rendering. In most cases, consistently oriented normal vectors associated to points are necessary, e.g., for surface reconstruction (Kazhdan et al., 2006; Kazhdan and Hoppe, 2013; Kazhdan et al., 2020; Lu et al., 2018). While locally estimating normals for a small patch of points can be done with simple techniques such as principal component analysis (PCA), obtaining a globally consistent orientation remains a notoriously difficult problem. Older methods mostly take a propagation-based approach by first estimating normals locally and then propagating the orientation globally (Hoppe et al., 1992; Xie et al., 2003; König and Gumhold, 2009; Seversky et al., 2011; Metzer et al., 2021). These methods are generally fast but less robust to noise, thin structures and sharp edges where propagation easily fails. Recent developments focus more on the global consistency by introducing new optimization formulations that involve the whole point cloud (Hou et al., 2022; Lin et al., 2022; Xu et al., 2023). Despite achieving the state-of-the-art performance, these global methods often suffer from high space/time complexity, preventing their practicability.
@@
An exciting and inspiring tread in this field is the exploration of various properties of the winding number formula (see Theorem 6). From the second line in Eq. (5) of the winding number formula, Lin et al. (2022) observed that the winding number field computed using correct normals should evaluate to $1/2$ at the input points. Lin et al. (2022) further turned this observation into a practical linear system, from which normals can be solved using conjugate gradients (Hestenes and Stiefel, 1952). Complementarily, Xu et al. (2023) use the first and the third line in Eq. (5) to define a regularization term such that the winding number evaluates to either 00 or $1$ almost everywhere. By turning the winding number formula to optimization objectives, these two methods above successfully achieve state-of-the-art performance in estimating consistently oriented normals. However, both methods suffer from high computational complexity. The linear system in PGR (Lin et al., 2022) has size $N\times N$ and is dense. Running PGR (Lin et al., 2022) with 40,000 points consumes about 8 GB GPU memory. The high memory usage prohibits it from scaling to large point clouds. On the other hand, GCNO (Xu et al., 2023) has high time complexity due to its objective function being non-linear. Solving a 10,000-point model using GCNO takes about one hour.
Figure 2. Illustration of the winding number normal consistency (WNNC) property. From left to right, we start with an input point cloud with normals, evaluate the winding number formula, and finally compute the negative field gradients at the input point positions. If the normals at the start are correctly oriented as shown in the first row, then the intermediate field is exactly the winding number field of the underlying shape. Consequently, the negative gradients at these point positions would be aligned with the input normals. We refer to this property as the winding number normal consistency (WNNC) property. Note that random normals do not satisfy the WNNC property (second row).
[…]
Illustration of the WNNC property.
Table 1. A brief comparison of time and memory consumption for the state-of-the-art methods GCNO (Xu et al., 2023), PGR (Lin et al., 2022), Dipole (Metzer et al., 2021) and iPSR (Hou et al., 2022). The results are obtained from running an Armadillo model (Krishnamurthy and Levoy, 1996) with uniform sampling on a desktop with an Intel i9-10900X CPU (3.70GHz, 10 cores) and an NVIDIA RTX 3090 GPU. Note that the cells colored in gray are estimates based on complexity, because GCNO and PGR are too costly to run on large-scale point clouds.
Method $10^{4}$ points $10^{5}$ points $10^{6}$ points
Time Mem. Time Mem. Time Mem.
GCNO (CPU) 1.1h ¡1G ¿1day 1.7G weeks ¿10G
PGR (GPU) 8s 1G ¿10min ¿30G ¿1day ¿3000G
Dipole (GPU) 17s 1G 90s 2.3G 230s 7.5G
iPSR (CPU) 18s ¡1G 50s ¡1G 380s 2.6G
Ours (CPU) 3s ¡1G 15s ¡1G 294s 1G
Ours (GPU) 2s ¡1G 6s ¡1G 86s 1.5G
@@
While it seems that PGR (Lin et al., 2022) and GCNO (Xu et al., 2023) have exhausted all possibilities with the winding number formula by using all three lines in Eq. (5) as constraints, we further exploit the winding number formula by exploring its higher-order properties. To be more specific, we propose to utilize the fact that, when using correctly oriented normals to evaluate the winding number formula, the normals (negative gradients) of the formula should be codirectional to the given normals. We refer to this property as the winding number normal consistency (WNNC). An illustration of WNNC is given in Fig. 2. While the WNNC seems to put a strong constraint on the possible orientations of normals, unfortunately, it cannot resolve the inside/outside orientation ambiguity. We thus propose to incorporate the objective function from PGR (Lin et al., 2022), which is known to encourage a consistently outward orientation but often converges at inaccurate normals. We propose to combine the advantages of both formulations by alternating between WNNC-based normal updates and PGR-based gradient descents, which leads to an embarrassingly simple yet effective iterative algorithm that allows fast and high-quality convergence to a globally consistent normal vector field. Furthermore, the proposed algorithm only involves repeatedly evaluating the winding number formula and its derivatives. Due to their special structures, these operations can be accelerated using a treecode-based algorithm (Barnes and Hut, 1986) and parallelized on GPU. Lacking an open-source implementation of the treecode algorithm that fully supports our needs (GPU-based and support for the winding number kernel derivative functions), we implement our custom treecode-based solver in CUDA. As shown in Table 1, our implementation significantly outperforms recent state-of-the-art methods in terms of space/time complexity. Moreover, our implementation is integrated with PyTorch (Paszke et al., 2019; Ansel et al., 2024) and will be made publicly available to facilitate future research into winding numbers.
@@
  • We implement a treecode-based acceleration algorithm in CUDA integrated with PyTorch, achieving substantial improvements over prior methods in terms of computational costs. We also make our code publicly available to facilitate further research.
2. Related Work
The research into estimating consistently oriented normals has a long history, dating back to decades ago. In this review, we mainly introduce methods that focus on the global consistency of normal orientation. Following prior work, we categorize existing methods into three types: propagation-based, volumetric and deep learning-based. We remark that these types are not completely mutually exclusive, and we introduce existing works based on their most prominent contributions.
2.1. Propagation-based Methods
@@
Local normal estimation is usually carried out by fitting a primitive geometry to local samples. Popular choices for this step include local tangent planes obtained from Principal Component Analysis (PCA) (Hoppe et al., 1992), implicit quadric surfaces (Xie et al., 2003, 2004), polynomials obtained by moving least squares (Levin, 2004), osculating jets (Cazals and Pouget, 2005; Metzer et al., 2021). These techniques are often efficient and well-established.
Once unoriented normals are available, it remains to determine a sign for each of them. The general practice is to propagate local orientations through neighbors to achieve global consistency. Hoppe et al. (1992) build a graph for input points whose edge costs are given by a measure of orientation consistency. Seeking a globally consistent orientation is then formulated as a graph optimization problem. Hoppe et al. (1992) use the minimum spanning tree (MST) as a greedy solution, which boils down to propagating the orientation along the MST. This simple greedy approach is efficient but lacks robustness. Many follow-up methods improve one or more aspects of Hoppe et al. (1992). To be robust to sharp features, Xie et al. (2003) start from multiple seed points with an improved flipping criterion to orient the whole point cloud. König and Gumhold (2009) further introduced Hermite curves into the flipping criterion to reduce the possibility of failures. Seversky et al. (2011) use harmonic functions on the point cloud to improve the constructed MST. Going beyond MST-based approaches, other propagation schemes have been proposed as well. Jakob et al. (2019) build a graph of points and equate a graph-edge collapse propagation with as an orientation step, and implement a fast parallel algorithm that collapse edges in a greedy manner. Instead of propagating orientations on the surface, Xie et al. (2004) first use locally oriented surfaces to determine locally oriented regions, and then propagate region orientations through a voting scheme. Recently, Metzer et al. (2021) propose to use neural networks to estimate consistently orientated normals for local patches, and then use dipole-based propagation to achieve global consistency.
@@
2.2. Volumetric Methods
Instead of directly estimating normals for point clouds, volumetric methods adopt an indirect approach where the whole volume is partitioned into inside/outside regions. Normal orientations can be determined as the direction pointing from inside to outside.
Early works make use of space partitioning structures to identify inside/outside regions. Dey and Goswami (2004) prove that, under certain assumptions, Delaunay balls of the Voronoi diagram can be well separated into outer balls and inner balls to give a desired partitioning. Chen et al. (2010) propose the binary orientation tree which gradually tags its node corners as inside/outside while it grows.
@@ 2.2. Volumetric Methods
For applying additional constraints, Alliez et al. (2007) make use of unoriented normals estimated from Voronoi cells, and solve for an implicit field that minimizes an anisotropic Dirichlet energy measuring the unoriented alignment between Voronoi-based normals and the implicit field gradient. Walder et al. (2005) apply value and gradient maximization together with smoothness regularization to an implicit function represented by radial basis functions (RBFs). Also based on RBFs, Huang et al. (2019) adopt triharmonic kernels and Duchon’s energy (Duchon, 1977) and achieve high robustness to point sparsity and non-uniform sampling.
@@
Signed distance fields (SDFs) are a popular choice of implicit fields. Mello et al. (2003) compute an approximate SDF on a tetrahedral grid. Zhao et al. (2001) solve the Eikonal equation to obtain the SDF on a rectangular grid. Recent methods also model SDFs as neural networks and apply extra regularization as loss functions, e.g., implicit geometric regularization with the Eikonal equation (Gropp et al., 2020), or enforcing the singularity of Hessian (Wang et al., 2023).
@@
Most closely related to our work are methods based on the winding number field (Barill et al., 2018; Lu et al., 2018). Simply speaking, the winding number field of an object equals $1$ on the inside and 00 on the outside. Since computing the winding number field requires known normals, several methods derive properties of the winding number field and use them in turn as constraints to obtain normals. Observing the fact that the winding number field is piece-wise constant (i.e., having zero Dirichlet energy), Takayama et al. (2014) solve for an orientation that minimizes the Dirichlet energy of the field computed with the winding number formula. Recently, Lin et al. (2022) derived a linear system formulation enforcing the field to equal $1/2$ on input samples, from which normals can be efficiently solved. Xu et al. (2023) designed a double-well loss function encouraging the field to evaluate to either 00 or $1$ almost everywhere. Concurrently, Liu et al. (2024) derive a boundary energy from the harmonicity of the winding number. Optimizing the boundary energy effectively leads to consistently oriented normals.
@@
Compared with propagation-based methods, most volumetric methods take a more global view, in the sense that their optimization objectives consider all input points throughout the whole execution process of the algorithm. As a result, these methods generally achieve much better global consistency but at a notably higher computational cost.
2.3. Deep Learning Methods
@@
In the past decade, deep learning has stormed over nearly all scientific fields, with normal estimation being no exception. The powerful fitting ability of neural networks allows direct inference of normals given an input point cloud. PCPNet (Guerrero et al., 2018) and Nesti-Net (Ben-Shabat et al., 2019) use neural networks to process multiscale point features and directly infer point normals. There are also volumetric learning-based methods that infer the implicit field rather than the normals (Erler et al., 2020; Xiao et al., 2022). Different from these straight-forward regression approaches, AdaFit (Zhu et al., 2021) predicts point offsets prior to normal estimation to reduce the number of outliers. NeAF (Li et al., 2023) further extends this idea by learning a neural angle field that offsets any given normal to the ground-truth one.
@@
2.4. Other Consistent Normal Estimation Methods
A few methods do not belong to the aforementioned categories. Shape as Points (SAP) (Peng et al., 2021) derives a differentiable version of Poisson Surface Reconstruction (PSR) (Kazhdan et al., 2006), and uses it to update normals such that the reconstructed mesh has a low error with the input points. Since the differentiable PSR (Peng et al., 2021) is performed in the frequency domain with fast Fourier transform, its reconstruction accuracy is limited by grid resolution. Iterative Poisson Surface Reconstruction (iPSR) (Hou et al., 2022) starts with applying Screened Poisson Surface Reconstruction (SPSR) (Kazhdan and Hoppe, 2013) to points with arbitrary normals, and iteratively resample normals from the mesh reconstructed in the previous iteration. Hou et al. (2022) show that this simple iterative procedure can update normals into a consistent orientation in only 5$\sim$30 of steps.
3. Preliminaries
Our method builds on the winding number formula (Barill et al., 2018) (also referred to as the Gauss formula in potential theory (Lu et al., 2018)), and is closely related to PGR (Lin et al., 2022). We begin with a brief review of them.
Theorem 3.1 (Winding Number).